# Hypencoder: Hypernetworks for Information Retrieval

### v1

@@ 4.4. Results and Discussion
Table 1. Comparison on in-domain evaluation datasets. The symbols next to each baseline indicate significance values with $p<0.05$. Note, that $\dagger$ is a group of baselines.
@@ 4.4.1. In-Domain Results
Hypencoder also beats both rerankers MonoBERT and cross-SimLM; demonstrating that reranking cannot make up for a weak retriever’s performance. Continuing in TREC-DL ’19 and ’20 we find that Hypencoder even surpassed RepLLaMA which is more than 60x larger and which also uses a significantly larger document embedding dimension of 4096. In fact the only model beating Hypencoder in nDCG@10 is ColBERTv2 which uses an embedding for every token in the document compared to Hypencoder’s fixed 768 dimension token. MSMARCO Dev results are also good with Hypencoder outperforming all the baselines and outperforming a few of the reference models such as SPLADE++ and MonoBERT.
Overall Hypencoder’s in-domain results are exceptionally strong given the simple training routine used, small encoder model size, and document representation size. To the best of our knowledge, Hypencoder sets a new record for combined TREC-DL ’19 and ’20 nDCG@10 with a 768 dimension dense document vector.
@@ 4.4.2. Out-of-Domain Results
Table 2. Out-of-domain results in nDCG@10. We only compare significance with BE-Base. Significance results with $p<0.05$ are shown with the $\clubsuit$ and $p<0.1$ are shown with $\diamondsuit$.
@@
Table 3. Evaluation metrics for the harder set of tasks which include TREC DL-HARD, TREC Tip-of-my-tongue TOT, and FollowIR. Significance is shown at $p<0.1$. For FollowIR we do not perform significance tests on BM25.
@@
Model nDCG@10 RR R@1000 nDCG@10 RR nDCG@1000 AP p-MRR nDCG@5 p-MRR AP p-MRR
@@ 4.4.3. Results on Harder Retrieval Tasks
The results on the harder retrieval tasks are in Table 3, like in the out-of-domain section we only consider the main baseline models and BM25. We can see that in the harder tasks Hypencoder remains dominant over the baseline models with higher retrieval metrics in all but one column. Additionally, the relative improvement compared to the in-domain results are higher (on all metrics that Hypencoder is the best for) suggesting that on harder tasks the added complexity that can be captured by Hypencoder’s similarity function is especially important. Additionally the high performance on TREC tip-of-the-tongue (TOT) and FollowIR indicate that Hypencoder adapts well to different domains through domain-specific fine-tuning.
@@
On the evaluated subset of TREC DL-HARD we see that Hypencoder has stronger precision metrics than the baselines by a large margin. As mentioned previously the higher relative improvement suggests that Hypencoder is especially dominant on harder tasks which, in part, explains its higher performance on TREC DL ’19 and ’20. Though on in-domain dataset Hypencoder does better or the same on recall metrics, on TREC DL-HARD BE-Base has higher recall than Hypencoder. We suspect that this may be because the relevance function that the q-net applies is not smooth, which has the benefit of being more discerning and likely accounts for some of the precision gains. However, if the q-net makes a mistake the non-smooth scoring could result in a much harsher score than the linear inner product is capable of producing.
@@
Moving to TREC tip-of-my-tongue (TOT) we see that Hypencoder continues to perform well. Tip-of-the-tongue is a complex retrieval task with long queries and passages and multiple aspects, the fact Hypencoder outperforms the baselines by a large margin validates the need for a more complex relevance function.
@@
Finally we have FollowIR which has three subsets – on all three Hypencoder has the best performance on the retrieval evaluation metrics of choice, in many cases by a sizable amount. Beside the retrieval evaluation metrics we also include p-MRR which is a metric released in the FollowIR (Weller et al., 2024a) paper. The metric measures the change in document ranks before and after an instruction is modified to see how well the model responses to the additional requirements. A p-MRR of 0 indicates no change in document rank based on the instruction change and a p-MRR of +100 indicates the documents were perfectly changed based on the instruction while -100 indicates the opposite. For additional details we refer readers to the original FollowIR paper (Weller et al., 2024a). As p-MRR is relative to each model’s performance before the instructions are modified it is not indicative of stand-alone retrieval performance. With that said, Hypencoder is the only model to achieve a positive p-MRR indicating it correctly modified the document ranking based on the instruction. This is no small feat as in the original FollowIR paper no retrieval model of the size of Hypencoder was able to get a positive p-MRR and even many much larger models trained on large instruction retrieval datasets could not get a positive result.
@@ 4.4.4. Analysis of Efficiency
To be able to search large-scale collections Hypencoder has to work well while only doing computation on a small subset of the document corpus. This led us to develop the efficient algorithm in Section 3.6. To quantify the performance we do a set of experiments varying the key parameters to see how each impacts both search quality and query latency. The results of these experiments can be seen in Figure 2. Note the document-to-document graph used has 100 neighbors per document. The figures show that each parameter has an important role in search quality. The largest role is played by $maxIter$ which can drastically reduce search performance if not set high enough, but which plateaus after around 12. The value of $nCandidates$ follows a similar pattern with a drastic increase followed by a plateau. The parameter $\tilde{C}$, the number of initial candidates, is unique in that it has a more gradual rise and is the only parameter that causes a decrease in effectiveness if raised too high. We suspect this happens because only $nCandidates$ are explored at each iteration but all candidates in $\tilde{C}$ are considered visited and thus are not explored in the future. This could mean that many good results from the initial candidates are not explored, leaving potentially good neighbors of these nodes undiscovered. Another interesting aspect of $\tilde{C}$ is that time decreases as it increases when early stopping is used. This is likely because high quality candidates are found sooner allowing search to end more quickly. In general, our approximation seems to behave as expected in terms of both time and retrieval performance.
@@
Search Type Query Lat. (ms) DL ’19 DL ’20
@@ 4.4.5. Impact of q-net Depth
Hypencoder performance suggests that having a more complex relevance function does indeed help improve retrieval performance as we had hypothesized. This raises the question: what is the optimal complexity of this relevance function. To answer this question we trained four versions of Hypencoder each trained to produce a q-net with a different number of layers. We selected [2, 4, 6, 8] as the number of q-net layers.
The results can be seen in Figure 3. The experiment shows that there is a benefit beyond two layers and that at least four is required for the best performance. Performance stays the same with six, indicating that this might be the point of diminishing returns. Lastly, eight layers decrease performance. There could be a number of reasons for this including that eight layers are harder to optimize or because effectively learning how to use all eight layers takes longer and thus might achieve better performance with extended training time.
@@
5. Conclusion
We propose a new class of retrieval model, the Hypencoder which overcomes the limitations of inner product based similarity functions that we prove to exist. Our model achieves a new state-of-the-art on TREC DL ’19 and ’20 for BERT sized encoder models with a single dense document vector and shows even stronger relative improvement on harder retrieval tasks such as tip-of-the-tongue queries. Further we demonstrate that learned relevance models can be applied to large-scale search corpus in an efficient way with our proposed approximate search algorithm. As Hypencoder is a flexible framework there is much interesting future work to explore, such as multi-vector document representations and corpus pretraining to name a few.
Acknowledgements.
This work was supported in part by the Center for Intelligent Information Retrieval, in part by the NSF Graduate Research Fellowships Program (GRFP) Award #1938059, and in part by the Office of Naval Research contract number N000142412612. Any opinions, findings and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect those of the sponsor.

### v2

@@ 4.4. Results and Discussion
Table 1. Comparison on in-domain evaluation datasets. The symbols next to each baseline indicate significance values with $p<0.05$. Note that $\dagger$ is a group of baselines.
@@ 4.4.1. In-Domain Results
Hypencoder also beats both rerankers MonoBERT and cross-SimLM; demonstrating that reranking cannot make up for a weak retriever’s performance. Continuing in TREC-DL ’19 and ’20 we find that Hypencoder even surpassed RepLLaMA which is more than 60x larger and which also uses a significantly larger document embedding dimension of 4096. In fact the only model beating Hypencoder in nDCG@10 is ColBERTv2 which uses an embedding for every token in the document compared to Hypencoder’s fixed 768 dimension embedding. MSMARCO Dev results are also good with Hypencoder outperforming all the baselines and outperforming a few of the reference models such as SPLADE++ and MonoBERT.
Overall, Hypencoder’s in-domain results are exceptionally strong given the simple training routine used, small encoder model size, and minimal document representation size. To the best of our knowledge, Hypencoder sets a new record for combined TREC-DL ’19 and ’20 nDCG@10 with a 768 dimension dense document vector.
@@ 4.4.2. Out-of-Domain Results
Table 2. Out-of-domain results in nDCG@10. We only compare significance with BE-Base. Significance results with $p<0.05$ are shown with $\clubsuit$ and $p<0.1$ are shown with $\diamondsuit$.
@@
Table 3. Evaluation metrics for the harder set of tasks which include TREC DL-HARD, TREC Tip-of-the-Tongue TOT, and FollowIR. Significance is shown at $p<0.1$.
@@
Model nDCG@10 MRR R@1000 nDCG@10 MRR nDCG@1000 AP p-MRR nDCG@5 p-MRR AP p-MRR
@@ 4.4.3. Results on Harder Retrieval Tasks
The results for the harder retrieval tasks are in Table 3, like in the out-of-domain section we only consider the main baseline models and BM25. We can see that in the harder tasks Hypencoder remains dominant over the baseline models with higher retrieval metrics in the vast majority of metrics. Additionally, the relative improvement is higher compared to the in-domain results (on all metrics where Hypencoder is the best) suggesting that on harder tasks the added complexity that can be captured by Hypencoder’s similarity function is especially important. Additionally, the high performance on TREC tip-of-the-tongue (TOT) and FollowIR indicate that Hypencoder adapts well to different domains through domain-specific fine-tuning.
@@
On the evaluated subset of TREC DL-HARD we see that Hypencoder has stronger precision metrics than the baselines by a large margin. As mentioned previously, the higher relative improvement suggests that Hypencoder is especially dominant on harder tasks which, in part, explains its higher performance on TREC DL ’19 and ’20. Though on in-domain dataset Hypencoder does better or the same on recall metrics, on TREC DL-HARD BE-Base has higher recall than Hypencoder. We suspect that this may be because the relevance function that the $q$-net applies is not smooth, which has the benefit of being more discerning and likely accounts for some of the precision gains. However, if the $q$-net makes a mistake the non-smooth scoring could result in a much harsher score than the linear inner product is capable of producing.
@@
Moving to TREC Tip-of-the-Tongue (TOT) we see that Hypencoder continues to perform well. Tip-of-the-tongue is a complex retrieval task with long multi-aspect queries and passages, the fact Hypencoder outperforms the baselines by a large margin validates the need for a more complex relevance function.
@@
Finally, we have FollowIR which has three subsets – on all three Hypencoder has the best performance on the retrieval evaluation metrics of choice, in many cases by a sizable margin. Beside the retrieval metrics, we also include p-MRR which is a metric introduced in the FollowIR (Weller et al., 2024a) paper. The metric measures the change in document ranks before and after an instruction is modified to see how well the model responds to the additional requirements. A p-MRR of 0 indicates no change based on the changed instruction and a p-MRR of +100 indicates the documents were perfectly changed while -100 indicates the opposite. As p-MRR is relative to each model’s performance before the instructions are modified it is not indicative of stand-alone retrieval performance. That being said, Hypencoder is the only model to achieve a positive p-MRR indicating it correctly modified the document ranking based on the instruction. This is no small feat as the FollowIR paper reports no retrieval model with the same size of Hypencoder was able to get a positive p-MRR and even many much larger models trained on large instruction retrieval datasets could not get a positive result.
@@ 4.4.4. Analysis of Efficiency
To be able to search large-scale collections Hypencoder has to work well while only doing computation on a small subset of the document corpus. This led us to develop the efficient algorithm in Section 3.6. To quantify the performance we do a set of experiments varying the key parameters to see how each impacts both search quality and query latency. The results of these experiments can be seen in Figure 2. Note that the document-to-document graph used has 100 neighbors per document. The figures show that each parameter has an important role in search quality. The largest role is played by $maxIter$ which can drastically reduce search performance if not set high enough, but which plateaus after around 12. The value of $nCandidates$ follows a similar pattern with a drastic increase followed by a plateau. The parameter $\tilde{C}$, the number of initial candidates, is unique in that it has a more gradual rise and is the only parameter that causes a decrease in effectiveness if raised too high. We suspect this happens because only $nCandidates$ are explored at each iteration but all candidates in $\tilde{C}$ are considered visited and thus are not explored in the future. This could mean that many good results from the initial candidates are not explored, leaving potentially good neighbors of these nodes undiscovered. Another interesting aspect of $\tilde{C}$ is that time decreases as it increases when early stopping is used. This is likely because high quality candidates are found sooner allowing search to end more quickly. In general, our approximation seems to behave as expected in terms of both time and retrieval performance.
@@
Search Type Query Latency (ms) DL ’19 DL ’20
@@ 4.4.5. Impact of $q$-net Depth
Hypencoder performance suggests that having a more complex relevance function does indeed help improve retrieval performance as we had hypothesized. This raises the question: what is the optimal complexity of this relevance function? To answer this question we trained four versions of Hypencoder each trained to produce a $q$-net with a different number of layers. We selected [2, 4, 6, 8] as the number of $q$-net layers.
The results in Figure 3 show that there is a benefit beyond two layers and that at least four are required for the best performance. Performance stays the same with six, indicating that this might be the point of diminishing returns. Lastly, eight layers decrease performance. There could be a number of reasons for this, including that eight layers are harder to optimize or because effectively learning how to use all eight layers takes longer and thus might achieve better performance with extended training time.
@@
5. Conclusions
We propose a new class of retrieval model—Hypencoder—that overcomes the limitations of similarity functions based on inner products that we prove to exist. Our model achieves a new state-of-the-art on TREC DL ’19 and ’20 for BERT sized encoder models with a single dense document vector and shows even stronger relative improvement on harder retrieval tasks such as tip-of-the-tongue queries. Additionally, we demonstrate that learned relevance models can be applied to large-scale search corpus in an efficient way with our proposed approximate search algorithm.
Acknowledgments. This work was supported in part by the Center for Intelligent Information Retrieval, in part by the NSF Graduate Research Fellowships Program Award #1938059, and in part by the Office of Naval Research contract number N000142412612. Any opinions, findings, conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect those of the sponsor.